The frequency of wave is 0.002 Hz. Its time period is
(a) 100 s (b) 500 s
(c) 5000 s (d) 50 s

Grade:Upto college level

1 Answers

Vikas TU
3 years ago
Dear student 
T = 1/f 
T = 1/0.002 = 1000/2 = 500 s 
Optin B is correct …...............................